# Binaural Target Speaker Extraction using HRTFs

### v3

@@ 4.2 Algorithm Setup
4.3 Metrics
To evaluate our method, we employed SI-SDR [27], wide-band and narrow-band Perceptual Evaluation of Speech Quality (PESQ[28], and Short-time Objective Intelligibility (STOI[29] as intrusive metrics of speech quality and intelligibility, i.e., measures that compare the extracted signal against a reference target. In addition, we used DNS-MOS [30], a non-intrusive DNN-based predictor of Mean Opinion Score (MOS), which estimates speech quality (SIG), background noise quality (BAK), and overall quality (OVRL).
[…]
@@ 4.3 Metrics
Fig. 3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced. $\tilde{\bm{s}}_{d}$ and $\tilde{\bm{s}}_{i}$ denotes the anechoic target signals. Graphs produced by [31].
Table 1: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples. Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap . Best scores are highlighted in bold.
@@
LS Unprocessed (2.688\3.380\3.140) -0.01 1.23 1.61 0.72 - -
BLCMV Beamformer [4] (2.827\3.304\3.495) -1.0 3.46 3.9 0.95 0.672 0.599
BDE-BiTSE [15] (3.072\3.471\3.768) 24.01 3.66 4.02 0.99 0.142 0.485
Bi-TSE-HRTF (proposed) (2.999\3.398\3.742) 20.18 3.14 3.78 0.98 0.1024 0.0042
MLS-French Unprocessed (2.499\3.252\2.947) -0.01 1.23 1.63 0.71 - -
BLCMV Beamformer (2.761\3.277\3.384) -1.03 3.37 3.82 0.94 0.694 0.608
BDE-BiTSE (2.987\3.403\3.709) 23.67 3.67 4.04 0.99 0.3 0.503
Bi-TSE-HRTF (proposed) (2.900\3.324\3.655) 19.86 3.14 3.79 0.98 0.1039 0.0041
@@ 4.4 Competing Methods
To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear). Accordingly, we compared our method against the BLCMV beamformer [4] and a recent DNN-based model [15], both specifically designed for this scenario. The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the same training data as ours. We ensured that the strongest variant reported in the paper (BDE+CDF+IPD+SDF) was used for comparison.
@@
In addition, we trained two variants of our model. The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are split into their RI components and concatenated, resulting in four input channels instead of two. The same decomposition is applied to the HRTF clue. The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue instead of the HRTF. In this case, the DOA is represented as a one-hot encoded vector with the same resolution as the available HRTFs, and passed through a fully connected layer to match the dimensionality of the bottleneck. For both variants, the loss function in (5) remains unchanged, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
4.5 Results
In Table 1, we present a comprehensive comparison of our method with competing algorithms across all datasets under both anechoic and reverberant conditions. In terms of SI-SDR, PESQ, and DNS-MOS, our approach achieves competitive results in the anechoic setting, though it underperforms in reverberant environments compared to BDE-BiTSE. Nevertheless, our model exhibits a clear advantage in preserving binaural cues, surpassing the baseline methods, as evident from the ITD and ILD measures.
Both DNN-based methods show strong cross-dataset generalization: though trained only on WSJ0-CSR1, they perform well on other corpora, including French, as shown in Table 1. We attribute this to the absence of speaker-specific priors and the reliance on a known HRTF database.
Additionally, two other variants of our model are evaluated, namely Bi-TSE-DOA and Bi-TSE-HRTF-RI. The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues, especially under reverberant conditions, underscoring the central role of HRTFs in our formulation. By contrast, the RI model achieves SI-SDR and binaural cue preservation comparable to the complex-valued model in the anechoic case, but suffers from reduced perceptual quality (lower PESQ) and fails to maintain cue preservation under reverberation. Figure 3 shows the joint ITDILD p.d.f.s across methods. The close alignment of our model’s distribution with that of the target underscores its effectiveness in maintaining spatial cues. In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations. The dereverberation effect is also apparent, as the joint ITD-ILD p.d.f. becomes more concentrated, indicating dominance of the direct path in the output of the proposed method. Overall, it is evident that the proposed approach achieves this more effectively than competing techniques. Table LABEL:tab:delta_thata further highlights the spatial consistency observed in Fig. 3. Specifically, we quantify the ITD error and the resulting DOA error, computed according to the procedure described in Sec. 4.3. The results clearly show that both competing methods exhibit substantial errors, whereas the proposed method preserves spatial consistency in the output signal.
@@ 4.5 Results
Fig. 4: A polar plot of SI-SDR versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue). Stars mark true DOAs; the dotted arc denotes the 0,dB threshold.
Finally, to quantify spatial resolution, we use a “scanning” procedure: signals are extracted with HRTFs at azimuths in $90^{\circ},90^{\circ}]$, and the SI-SDR of each is computed against the target. Plotting these values yields a beam-pattern-like diagram (Fig. 4) with a clear peak at the true DOA. The main lobe width characterizes the model’s stability when the HRTF clue deviates mildly from the ground truth.
5 Conclusions
@@
This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, yielding a personalized solution to the task at hand. Through cross-dataset evaluations, we demonstrate robust generalization across diverse corpora. The proposed approach not only produces high-quality speech but also faithfully preserves the binaural characteristics of the signal. Comparative analyses with existing binaural TSE methods reveal substantial improvements in spatial consistency, while preserving high speech quality. These findings highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction under a wide range of acoustic conditions.

### v4

@@ IV-B Algorithm Setup
Figure 3: The joint p.d.f.s of ITD [ms] and ILD [dB] for $\theta_{d}=-60^{\circ}$, $\theta_{i}=-90^{\circ}$, and $\phi_{d,i}=-10^{\circ}$, with $T_{60}=0.63$ s, for the frequency band centerd at 500 Hz, where the ILD is less pronounced. $\tilde{\bm{s}}_{d}$ and $\tilde{\bm{s}}_{i}$ denotes the anechoic target signals. Graphs produced by [27].
IV-C Metrics
To evaluate our method, we use SI-SDR [28], wide- and narrow-band Perceptual Evaluation of Speech Quality (PESQ[29], and Short-time Objective Intelligibility (STOI[30] as intrusive metrics of speech quality and intelligibility. In addition, we employ DNS-MOS [31], a non-intrusive DNN-based Mean Opinion Score (MOS) predictor estimating speech quality (SIG), background noise quality (BAK), and overall quality (OVRL). Furthermore, to assess binaural cue preservation, we compare the ILDITD p.d.f.s of the extracted and target signals, where close agreement indicates preserved spatial cues. These metrics are computed using the procedure in [27].
TABLE I: Each evaluation dataset contains 1500 speech mixtures, with both speakers extracted, yielding 3000 samples. Mixtures were created with SIR drawn from $\mathcal{U}[-5,5]$ dB, and 100% overlap. Best scores are highlighted in boldface.
@@ IV-D Competing Methods
To ensure a fair comparison, all baseline methods were evaluated under an identical setup, namely binaural TSE with only two microphones (one per ear). Accordingly, we compared our method against the BLCMV beamformer [4] and a recent DNN-based model [15], both specifically designed for this scenario. The competing DNN, referred to as BDE-BiTSE, was trained following the procedure described in the paper, using the exact same training data as ours. We ensured that the strongest variant reported in the paper (BDE+CDF+IPD+SDF) was used for comparison.
@@
As an ablation study, we trained two variants of our model. The first, Bi-TSE-HRTF-RI, is a real-valued model in which the complex-valued input features are decomposed into their RI components and concatenated, yielding four input channels rather than two. The same decomposition is applied to the HRTF clue. The second, Bi-TSE-DOA, is a complex-valued model that incorporates the DOA as a clue rather than the HRTF. Here, the DOA is represented as a one-hot-encoded vector with the same resolution as the available HRTFs and passed through a fully connected layer to match the bottleneck’s dimensionality. For both variants, the loss function (5) remains unaltered, with the target defined as the binaural rendering of the desired speech signal using the correct HRTF.
IV-E Results
[…]
Figure 3 shows the joint ITDILD p.d.f.s for the extracted and target signals. The close alignment of the proposed method with the target distribution confirms effective preservation of binaural cues. In contrast, competing methods exhibit noticeable deviations, resulting in shifted perceived source locations. The increased concentration of the joint distribution further indicates a dereverberation effect, with the direct path dominating the output.
@@ IV-E Results
Figure 4: A polar plot of SI-SDR [dB] versus DOA, with the target at $\theta_{d}=-75^{\circ}$ (red) and the interferer at $\theta_{i}=-10^{\circ}$ (dashed blue). Stars mark true DOAs;
Additionally, we evaluate two model variants: Bi-TSE-DOA and Bi-TSE-HRTF-RI. The DOA-based model yields lower SI-SDR and PESQ, fails to suppress the interferer, and does not preserve binaural cues—especially under reverberation—highlighting the central role of HRTFs in our formulation. The RI model achieves SI-SDR and binaural cue preservation comparable to those of the complex-valued model in anechoic conditions, but yields lower PESQ scores and degraded cue preservation under reverberation. To demonstrate spatial resolution, we employ a “scanning” procedure in which signals are extracted using HRTFs over azimuths in $90^{\circ},90^{\circ}]$, and the corresponding SI-SDR is computed with respect to the target. Plotting these values yields a beam-pattern-like diagram (Fig. 4) with a peak at the true DOA, where the main-lobe width reflects the model’s stability to mild HRTF deviations.
V Conclusions
@@
This paper presents a speaker-independent TSE method that exploits the listener’s HRTFs in both anechoic and reverberant environments, providing a personalized solution. By conditioning the extraction process on individualized HRTFs, the proposed approach preserves binaural structure while maintaining speech quality. Comparative analyses with existing binaural TSE methods demonstrate substantial improvements in spatial consistency, while preserving high speech quality. These results highlight the critical role of the listener’s HRTF in anchoring spatial cues, enabling accurate and consistent target-speaker extraction across diverse acoustic conditions.